OpenStudy (anonymous):

If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. Determine the original fraction. Which of the following equations represents "If the numerator of a fraction is increased by 3, the fraction becomes 3/4"? (Hint: cross products)

OpenStudy (vivek3461):

let the required fraction be \[\frac{ x }{ x+7 }\] if the numerator is increased by 3, then \[\frac{ x+3 }{ x+7 }=\frac{ 3 }{ 4 }\] by cross multiplication we get, 4 ( x +3) = 3 (x+7) 4x+12=3x+21 4x-3x=21-12 x=9 So the required fraction is 9 / 16